OpenStudy (anonymous):

Use the distance formula to fond the distance between the points (1,-3) and -5,6)

OpenStudy (anonymous):

\[\sqrt{(1+5)^{2}+(-3-6)^{2}}\]

OpenStudy (anonymous):

underroot(81+36)=(117)^1/2

OpenStudy (anonymous):

I'm not sure i get the same the way I did it. A. 6.71 b. 45 c. 10.82 d. 10 e. 73

OpenStudy (anonymous):

I got square root of 51 giving 7.14... Etc

OpenStudy (anonymous):

But it's not one of the options

OpenStudy (anonymous):

10.82 is correct ans

OpenStudy (anonymous):

Let d be the distance. d=√((〖(-5-1)〗^2+〖(6+3)〗^2 ) ) . d=√(36+81 ) =√117